OpenStudy (anonymous):

MEDEAL AND FAN Find the sum or difference. 1. Mr.Jensen bought 2 2/3 gallons of gasoline yesterday. Today he bought 3 1/2 more gallons. How many total gallons of gasoline has he bought? A 5 3/5 gallons B 5 3/6 gallons C 6 1/6 gallons D 5 1/6 gallons

OpenStudy (solomonzelman):

do you know how to do these type of problems? lets say it was: ``` Sam bought 4 apples yesterday, and today he bought another 3 apples. How many apples does Sam now have? ``` then you would add the apples to get your answer. 3+4=7 So the answer is 7.

OpenStudy (solomonzelman):

your example is not any different, besides that it involves adding fractions.

OpenStudy (anonymous):

\[2\tfrac{2}{3}+3\tfrac{1}{2}\] is what you have to add the \(2+3=5\) part is pretty easy, all that is left is \(\frac{2}{3}+\frac{1}{2}\)